dt0123:

Find the value of tan(3Arc cos3/4)

sillybilly123:

\(\arccos \frac{3}{4}\) is an angle, so let \(\theta = \arccos \frac{3}{4} \implies \cos \theta = \frac{3}{4}\) so hypoteneous = 4, adjacent side = 3 means opposite side, from Pythagoreas, \(= \sqrt{7}\) Hence \(\tan \theta = \dfrac{\sqrt7}{3}\) Next use this: \( \tan 3\theta = \dfrac{3 \tan\theta - \tan^3\theta}{1 - 3 \tan^2\theta} \) \( = \dfrac{3 \frac{\sqrt7}{3} - (\frac{\sqrt7}{3})^3}{1 - 3 (\frac{\sqrt7}{3})^2} \) \(= -\dfrac{5 \sqrt(7)}{ 9}\) \(tan 3 \theta\) can be found using tan double angle formula